# CogniSNN: A First Exploration to Random Graph Architecture based Spiking Neural Networks with Enhanced Expandability and Neuroplasticity

### v1

@@ 4.1.1 ResNode
$\begin{split}O_{i}^{1}[t]&=ConvBNSN_{i}^{1}(I_{i}[t])\\[10.0pt] O_{i}^{2}[t]&=ConvBNSN_{i}^{2}(O_{i}^{1}[t])\\[10.0pt] O_{i}[t]&=OR(O_{i}^{2}[t],O_{i}^{1}[t])\\ &=O_{i}^{2}[t]+O_{i}^{1}[t]-O_{i}^{2}[t]\odot O_{i}^{1}[t]\end{split}$ (13)
In particular, $O_{i}^{1}[t]$ is derived from $ConvBNSN_{i}^{1}$ and is regarded as an identity. It is logically integrated with $O_{i}^{2}[t]$, which is obtained from $ConvBNSN_{i}^{2}$. Instead of an ADD gate, an OR gate ($OR(\cdot)$) is used because it enables the output of node $O_{i}[t]$ to maintain a spiking format rather than containing real values. Subsequently, $O_{i}[t]$ is transmitted to the descendant nodes of $v_{i}$. Residual learning facilitates the training of deep neural networks through identity mapping He et al. (2016b). Similarly, we have the following theorem.
@@
Due to limited space, the proof of Corollary 1 and the illustration of Property 1 are given in Section 3 of the Supplementary Material.
@@ 4.1.2 Tailored Pooling Mechanism
Given a node $v_{j}\in\mathcal{V}$ (not the root node), we estimate $D(I_{j}[t])$ according to the outputs from its predecessors $\mathcal{P}_{j}$ as follows.
@@ 4.2 Critical Path-Based LWF
To capture the neuroplasticity of CogniSNN as elucidated in the Introduction, we focus on parameter fine-tuning associated with the selective substructures, termed critical paths to achieve the neural reuse to reshape functions in response to new experiences. This idea is inspired by visualizing the activation frequency of nodes. As shown in Figure 11 of Supplementary Material, it can be observed that specific ResNodes and certain paths (defined as sequences of nodes in CogniSNN with information flow) are activated more/less frequently than others. Based on such observation, a critical path is targeted based on the concept of betweenness centrality from graph theory which is usually used to quantify the importance of a node (or edge) based on its role in connecting other nodes. In this paper, given a path $p$ formulated as:
@@
where $l_{p}$ is length of $p$, indicating the number of edges involved in $p$, we define the betweenness centrality of $p$, denoted by $C_{B}(p)$ as follows.
@@
In Definition 1, $C_{B}(v_{i})$ and $C_{B}(e_{j})$ are the betweenness centrality of a node and an edge, which follows the principle of graph theory  [^1] The definition of node betweenness centrality and edge betweenness centrality are in Section 4 of Supplementart Material.. Based on the definition of $C_{B}(p)$, we sort all paths in descending order according to their betweenness centrality:
@@
where $C_{B}(p_{1})\geq C_{B}(p_{2})...\geq C_{B}(p_{|P|})$, $|P|$ is the total number of paths. We then define the critical paths ($\mathcal{C}$) as a set of paths with capacity $K$ according to the definition of $C_{B}(p)$ and the concept related to the similarity between old and new tasks, mathematically formulated as follows.
@@
$\mathcal{C}\triangleq\{C_{B}(p_{k})|k\in[1,K],K\leq|P|\}$ if two tasks are similar. Otherwise, $\mathcal{C}\triangleq\{C_{B}(p_{k})|k\in[|P|-K+1,|P|],K\leq|P|\}.$
To measure the similarity between the datasets, we use the Fréchet Inception Distance (FID) Heusel et al. (2017), which is originally designed to evaluate generative models by comparing the differences in the distribution of the generated and real images in the feature space. Based on empirical observations, we consider two datasets similar when their FID is below the threshold of 50. For example, CIFAR100 and CIFAR10 show strong similarity with an FID of 24.5, while CIFAR100 and MNIST, having an FID of 2341.2, show weak similarity. This definition aligns with human cognitive processes, where principal neural pathways are leveraged for similar tasks, while unused connections are activated for unfamiliar tasks to assimilate new information and avoid forgetting prior knowledge. For simplicity, we set $K=1$ in this study.
@@
We design a continual learning algorithm implemented on CogniSNN by integrating critical paths and LWF method outlined in Section 3.2. The details are given in Algorithm  1. More precisely, $X^{+}$ and $Y^{+}$ are the dataset and label set in the new task. $CogniSNN$ is the model pre-trained based on the dataset in the old task. $\Theta$ and $\Theta^{+}$ are classifier layers in $CogniSNN$ and $CogniSNN^{+}$ where $CogniSNN^{+}$ is the retrained $CogniSNN$ based on $(X^{+},Y^{+})$. In line 2, we align the dimensions of $\Theta$ and $\Theta^{+}$ by adding or removing neurons in $\Theta$. In line 3, we freeze all parameters except $\theta_{\mathcal{C}}$ and $\theta_{\Theta^{+}}$ that are associated with $\mathcal{C}$ and $\Theta^{+}$. From lines 4 to 8, LWF is used to refine $\theta_{\mathcal{C}}$ and $\theta_{\Theta^{+}}$ with the dataset in the new task.
Theoretically, this critical path can be combined with many basic continual learning methods. In this paper, we only combine the more biologically plausible LWF approach for initial validation and exploration. In the future, we will combine more continual learning methods to explore the critical path more richly.
@@ 5.1 Experimental Setting
DVS-Gesture Amir et al. (2017) SGLFormer Zhang et al. (2024) SGLFormer-3-256 2.17 16 98.6
Spikformer Zhou et al. (2022) Spikformer-2-256 2.57 5 79.5
SSNN Ding et al. (2024) VGG-9 - 5 / 8 90.7 / 94.91
CogniSNN(Ours) ER-RAG-7 0.13 5/8/16 94.8/ 95.8/ 98.6
CogniSNN(Ours) WS-RAG-7 0.13 5/8/16 95.5/ 96.5/ 97.2
CIFAR10-DVS Li et al. (2017) Spikformer Zhou et al. (2022) Spikformer-2-256 2.57 5 68.6
STSA Wang et al. (2023) STSFormer-2-256 1.99 16 79.9
SSNN Ding et al. (2024) VGG-9 - 5 / 8 73.6 / 78.6
CogniSNN(Ours) ER-RAG-7 1.51 5/8/16 75.8/ 75.9/ 76.7
CogniSNN(Ours) WS-RAG-7 1.51 5/8/16 75.4/ 74.9/ 76.4
N-Caltech101 Li et al. (2017) MLFFeng et al. (2022) VGG-9 - 5 70.4
Spikformer Zhou et al. (2022) Spikformer 2.57 5 72.8
SSNN Ding et al. (2024) VGG-9 - 5/8 77.9/ 79.3
CogniSNN(Ours) ER-RAG-7 1.51 5/8 80.6/ 79.3
CogniSNN(Ours) WS-RAG-7 1.51 5/8 78.6/ 80.2
Table 1: Comparsion with other benchmark results on DVS-Gesture, CIFAR10-DVS, N-Caltech101.
We conduct experiments on three neuromorphic datasets on the NVIDIA RTX 4090 GPU. Furthermore, to demonstrate the advanced performance of CogniSNN, we engage in a detailed comparative evaluation against the SOTA SNN models, which are distinguished by architectures such as chain-like connectivity, transformer-based, and ResNet-based designs. To save space, more details of the datasets and configuration of the hyperparameters are provided in Section 1 of Supplemental Material.
@@ 5.2 Preliminary Analysis
Figure 2: Comparison of results on three neuromorphic datasets between CogniSNN and chain-like structured SNN.
Before deep analysis of the performance of CogniSNN, we first assess the feasibility and potential continual learning capacity of RGA-based SNN by comparing CogniNN with a chain-structured SNN. Both are arranged with an identical number of ResNodes. This preliminary analysis aims to motivate an in-depth exploration of RGA-based SNN modeling.
As for the potential of expandability, CogniSNN configured with 7 ResNodes consistently outperforms chain-structured SNNs, as seen in Figure 2. Random node connections act as implicit skip connections, dynamically prioritizing features by adjusting weights. This overcomes the limitation of sequentially increasing the dimensions of features with layers in traditional neural networks.
Model CIFAR100 CIFAR10 CIAFR100
@@
Table 2: Neuroplasticity for RGA-based SNN in similar tasks.
@@
Table 3: Neuroplasticity for RGA-based SNN in dissimilar tasks.
Regarding neuroplasticity, we compare CogniSNN configured by 32 ResNodes with chain-like structure-based SNN under the implementation of the proposed critical path-based LWF algorithm with $K=1$. Table 2 and 3 show the performance of the critical path-based LWF applied to similar tasks (FID = 24.5) and dissimilar tasks (FID = 2341.2), respectively. In particular, CIFAR 100 is treated as source data and CIFAR10 and MNIST are treated as target data. The last column indicates the classification accuracy in the source data, serving as a benchmark for comparison. Our CogniSNN exhibits exceptional classification capabilities for new tasks, achieving a minimal accuracy decline, notably a ${}48\%$ improvement in the MNIST dataset over chain-like models and a ${}18\%$ increase relative to the benchmark. This illustrates the feasibility of RGA-based SNNs in the realm of continual learning.
@@ 5.3 Insightful Performance Analysis
Table 1 presents the results of the experiments conducted with CogniSNN and comparative models. Two distinct CogniSNN architectures based on ER and WS are developed, respectively. The visualization of these architectures is provided in Section 2 of the Supplementary Material. Clearly, when tested with the time step $T=5$, our methodology outperforms the SSNN model by $5\%$ on DVS-Gesture, $2.2\%$ on CIFAR10-DVS and $2.6\%$ on N-Caltech101.
However, at $T=8$ and $T=16$, it should be noted that the performance on the CIFAR10-DVS is marginally inferior. In this regard, we believe that the graph structure is an important factor for the model performance. In future work, we will explore how to optimize the graph structure in order to achieve a better performance.
With $T=16$, it is obvious that transformer-based SNNs (e.g., SCLFormer and STSA) demonstrate strong performance in datasets such as DVS-Gesture and CIFAR10-DVS. However, our model achieves comparable accuracy while utilizing significantly fewer parameters. For example, when evaluated on DVS-Gesture, our model requires only 1/20 of the parameters needed by the transformer-based model (0.13 million versus 2.17 million).
@@
Figure 3: Comparison of results of OR and other operations on DVS-Gesture with ER-RAG-7 based CogniSNN.
5.4 Expandability
The expandability of CogniSNN is heavily based on the usage of skip connections within ResNodes. As described in Section 2.1, early research is generally based on the following four configurations: 1) without skip connection, 2) ADD (addition), 3) AND, and 4) IAND. We compare our skip connection (OR gate) with the four aforementioned configurations based on the ER-driven CogniSNN. Furthermore, all models with/without configuring different skip connection mechanisms are trained on the DVS-Gesture.
@@ 5.4 Expandability
5.5 Neuroplasticity
We apply the critical path-based LWF with $K=1$ on the WS-driven CogniSNN configured with 32 ResNodes. Comparable results are observed with the ER-driven CogniSNN and are provided in Section 6 of Supplementary Material. Tables 4 and 5 present the results, maintaining the significance of each column consistent with the corresponding column in Tables 2 and 3. The first row presents results from the application of the fine-tuned CogniSNN using traditional LWF on both tasks. The subsequent two rows display results utilizing the proposed critical path-based LWF, implemented based on paths with the highest and lowest betweenness centrality, respectively.
In Table 4, the performance of CogniSNN with fine-tuned parameters associated with the path in high $C_{B}$ is marginally more effective than fine-tuned based on low $C_{B}$, showing an improvement in $0.4\%$. Similarly, the performance of low $C_{B}$ improves $0.2\%$ than that of high $C_{B}$ in Table 5. The improvement is minor due to the simplicity of the classification task in MINST and CIFAR10. However, this supports our hypothesis regarding the correlation between the selection of the critical path and the similarity of the task (Definition 2).
Relative to the benchmark, forgetting occurs. Yet, when compared specifically to the LWF-based method, our algorithm demonstrates a $3.6\%$ and $5.1\%$ reduction in forgetting on the previous task. This confirms that the critical path-based LWF can effectively mitigate catastrophic forgetting.
Method CIFAR100 CIFAR10 CIAFR100

### v2

@@ 4.1.1 ResNode
$\begin{split}O_{i}^{1}[t]&=ConvBNSN_{i}^{1}(I_{i}[t])\\[10.0pt] O_{i}^{2}[t]&=ConvBNSN_{i}^{2}(O_{i}^{1}[t])\\[10.0pt] O_{i}[t]&=OR(O_{i}^{2}[t],O_{i}^{1}[t])\\ &=O_{i}^{2}[t]+O_{i}^{1}[t]-O_{i}^{2}[t]\odot O_{i}^{1}[t]\end{split}.$ (13)
In particular, $O_{i}^{1}[t]$ is derived from $ConvBNSN_{i}^{1}$ and is regarded as an identity. It is logically integrated with $O_{i}^{2}[t]$, which is obtained from $ConvBNSN_{i}^{2}$. Instead of an ADD gate, an OR gate ($OR(\cdot)$) is used because it enables the output of node $O_{i}[t]$ to maintain a spiking format rather than containing real values. Subsequently, $O_{i}[t]$ is transmitted to the descendant nodes of $v_{i}$. Residual learning facilitates the training of deep neural networks through identity mapping [2016_he_identity]. Similarly, we state the following theorem.
@@
In the ResNode, when implementing identity mapping, this residual structure can overcome gradient vanishing/explosion. With $O_{i}^{2}\equiv 0$, the gradient of the output of ResNode relative to the input of the residual block $ConvBNSN_{i}^{2}$ (the output of $ConvBNSN_{i}^{1}$) can be calculated as:
$\frac{\partial O_{i}[t]}{\partial O_{i}^{1}[t]}=\frac{\partial OR(O_{i}^{2}[t],O_{i}^{1}[t])}{\partial O_{i}^{1}[t]}=\frac{\partial O_{i}^{1}[t]}{\partial O_{i}^{1}[t]}=1.$ (14)
We do not consider the $ConvBNSN_{i}^{1}$ when calculating the gradient, because the input of $ConvBNSN_{i}^{1}$ is weighted and summed by multiple predecessors. Thus, we allow some loss in this part. This illustrates why the OR gate can overcome network degradation.
@@ 4.1.2 Tailored Pooling Mechanism
Given a node $v_{j}\in\mathcal{V}$ (not the root node), we estimate $D(I_{j}[t])$ according to the outputs from its predecessors $\mathcal{P}_{j}$ as follows:
@@ 4.2 Critical Path-Based LwF
To capture the path-plasticity of CogniSNN as elucidated in the Introduction, we focus on parameter fine-tuning associated with the selective substructures, termed critical paths to achieve the neural reuse to reshape functions in response to new experiences. This idea is inspired by visualizing the activation frequency of nodes. As shown in Figure 9 of Supplementary Material  [^1] The supplementary material and source code are both available at https://github.com/Yongsheng124/CogniSNN., it can be observed that specific ResNodes and certain paths (defined as sequences of nodes in CogniSNN with information flow) are activated more/less frequently than others. Based on such observation, a critical path is targeted based on the concept of betweenness centrality from graph theory, which is commonly used to quantify the importance of a node (or edge) based on its role in connecting other nodes. In this paper, given a path $p$ formulated as:
@@
where $l_{p}$ is the length of $p$, indicating the number of edges involved in $p$.
Then we define the betweenness centrality of $p$, denoted by $C_{B}(p)$ as follows.
@@
In Definition 1, $C_{B}(v_{i})$ and $C_{B}(e_{j})$ are the betweenness centrality of a node and an edge, which follow the principle of graph theory  [^2] The definition of node betweenness centrality and edge betweenness centrality are in Section 3 of Supplementary Material.. Based on the definition of $C_{B}(p)$, we sort all paths in descending order according to their betweenness centrality:
@@
where $C_{B}(p_{1})\geq C_{B}(p_{2})...\geq C_{B}(p_{|P|})$, $|P|$ is the total number of paths.
We then define the critical paths ($\mathcal{C}$) as a set of paths with capacity $K$ according to the definition of $C_{B}(p)$ and the concept related to the similarity between old and new tasks, mathematically formulated as follows.
@@
$\mathcal{C}\triangleq\{p_{k}|k\in[1,K],K\leq|P|\}$ if two tasks are similar. Otherwise, $\mathcal{C}\triangleq\{p_{k}|k\in[|P|-K+1,|P|],K\leq|P|\}$.
This definition aligns with human cognitive processes, where principal neural pathways are leveraged for similar tasks, while unused connections are activated for unfamiliar tasks to assimilate new information and avoid forgetting prior knowledge. For simplicity, we set $K=1$ in this study.
@@
We design a continual learning algorithm implemented on CogniSNN by integrating critical paths and the LwF method outlined in Section 3.2. The details are given in Algorithm  1. More precisely, $X^{+}$ and $Y^{+}$ are the dataset and label set in the new task. $CogniSNN$ is the model pre-trained based on the dataset in the old task. $\Theta$ and $\Theta^{+}$ are classifier layers in $CogniSNN$ and $CogniSNN^{+}$, respectively. Here, $CogniSNN^{+}$ denotes the retrained $CogniSNN$ based on $(X^{+},Y^{+})$. $\mathcal{I}$ represents the number of learning rounds for new model. In line 2, we align the dimensions of $\Theta$ and $\Theta^{+}$ by adding neurons in $\Theta$. In line 3, we freeze all parameters except $\theta_{\mathcal{C}}$ and $\theta_{\Theta^{+}}$ that are associated with $\mathcal{C}$ and $\Theta^{+}$. From lines 4 to 8, LwF is used to refine $\theta_{\mathcal{C}}$ and $\theta_{\Theta^{+}}$ with the data in the new task.
To measure the similarity between the datasets, we use the Fréchet Inception Distance (FID) [2017_heusel_gans], which is originally designed to evaluate generative models by comparing the differences in the distribution of the generated and real images in the feature space. Based on empirical observations, we consider two datasets similar when their FID is below the threshold of 50. For example, CIFAR100 and CIFAR10 show strong similarity with an FID of 24.5, while CIFAR100 and MNIST, having an FID of 2341.2, show weak similarity.
Theoretically, this critical path can be combined with many basic continual learning methods. In this paper, we only apply the more biologically plausible LwF approach for initial validation and exploration. In the future, we will integrate more continual learning methods to explore the critical path more richly.
@@ 5.1 Experimental Setting
Table 1: Comparison with other benchmark results on DVS-Gesture, CIFAR10-DVS, N-Caltech101 and Tiny-ImageNet.
Dataset Method Network Param(M) T Accuracy($\%$)
DVS-Gesture [2017_amir_low] MLF [2022_feng_multi] VGG-9 - 5 85.77
SEW-ResNet [2021_fang_deep] SEW-ResNet 0.13 16 97.90
SGLFormer [2024_zhang_sglformer] SGLFormer-3-256 2.17 16 98.60
CML [2023_zhou_enhancing] Spikformer-4-384 2.57 16 98.60
Spikformer [2022_zhou_spikformer] Spikformer-2-256 2.57 5 / 16 79.52 / 98.30
SSNN [2024_ding_shrinking] VGG-9 - 5 / 8 90.74 / 94.91
@@
[…]
CIFAR10-DVS [2017_li_cifar10] Spikformer [2022_zhou_spikformer] Spikformer-2-256 2.57 5 68.55
MLF [2022_feng_multi] VGG-9 - 5 67.07
AutoSNN [2024_shi_towards] AutoSNN - 8 72.50
SEW-ResNet [2021_fang_deep] SEW-ResNet 1.19 8 70.20
SSNN [2024_ding_shrinking] VGG-9 - 5 / 8 73.63 / 78.57
[…]
N-Caltech101 [2017_li_cifar10] MLF [2022_feng_multi] VGG-9 - 5 70.4
Spikformer [2022_zhou_spikformer] Spikformer 2.57 5 72.8
EventMix [2023_shen_eventmix] ResNet-18 - 10 79.5
TIM [2024_shen_tim] Spikformer - 10 79.0
SSNN [2024_ding_shrinking] VGG-9 - 5 / 8 77.97/ 79.25
[…]
Tiny-ImageNet [2015_le_tiny] Joint-SNN [2023_guo_joint] VGG-16 - 4 55.39
SNASNet-Bw [2022_kim_neural] Searched - 5 54.60
Online LTL [2022_yang_training] VGG-13 - 6 55.37
[…]
We conduct experiments five times on three neuromorphic datasets and one static dataset on the NVIDIA RTX 4090 GPU by SpikingJelly [2023_fang_spikingjelly]. Furthermore, to demonstrate the advanced performance of CogniSNN, we engage in a detailed comparative evaluation against the SOTA SNN models, which are distinguished by architectures such as chain-like connectivity, transformer-based, and ResNet-based designs. More details of the datasets and the hyperparameter configuration are provided in Section 1 of Supplemental Material.
@@ 5.2 Preliminary Analysis
Before deep analysis of the performance of CogniSNN, we first assess the feasibility and potential continual learning capacity of RGA-based SNNs by comparing CogniSNN with a chain-structured SNN. Both are arranged with an identical number of ResNodes. This preliminary analysis aims to motivate an in-depth exploration of RGA-based SNN modeling.
As for the potential of depth-scalability, CogniSNN configured with 7 ResNodes consistently outperforms chain-structured SNNs, as seen in Figure 2. Random connections among ResNodes act as implicit skip connections, dynamically prioritizing features by adjusting weights. This overcomes the limitation of sequentially increasing the dimensions of features with layers in traditional neural networks.
Figure 2: Comparison of results on three neuromorphic datasets between CogniSNN and the chain-like structured SNN.
Regarding path-plasticity, under the implementation of the vanilla LwF algorithm, we compare CogniSNN with 32 ResNodes against a chain-like structure-based SNN. Table 2 and 3 show the potential of CogniSNN for continual learning in similar tasks and dissimilar tasks, respectively. In particular, CIFAR100 is treated as source data and CIFAR10 and MNIST are treated as target data. The last column indicates the classification accuracy in the source data, serving as a benchmark for comparison. Our CogniSNN exhibits exceptional classification capabilities for new tasks, achieving a minimal accuracy decline, notably a $48\%$ improvement in the MNIST dataset over the chain-like model and a $18\%$ increase relative to the benchmark. This illustrates the feasibility of RGA-based SNNs in the realm of continual learning.
Table 2: Path-plasticity’s potential for CogniSNN in similar tasks.
Model CIFAR100 CIFAR10 CIFAR100
@@
Table 3: Path-plasticity’s potential for CogniSNN in dissimilar tasks.
@@ 5.3 Insightful Performance Analysis
Table 1 presents the results of the experiments conducted with CogniSNN and comparative models. Clearly, when tested with the time step $T=5$, our methodology outperforms the SSNN model by $4.04\%$ on DVS-Gesture, $2.19\%$ on CIFAR10-DVS and $2.67\%$ on N-Caltech101.
However, at $T=8$, it should be noted that the performance on the CIFAR10-DVS is marginally inferior. In this regard, we believe that the vanilla convolution module in ResNode is an important factor for the model performance. In future work, we will explore more advanced modules like self-attention [2017_vaswani_attention] to replace the convolution here, which we believe can achieve better performance.
With $T=16$, transformer-based SNNs (e.g., CML) demonstrate strong performance in datasets such as DVS-Gesture. However, our model achieves comparable accuracy while utilizing significantly fewer parameters. For example, when evaluated on DVS-Gesture, our model requires only 1/20 of the parameters needed by the CML model (0.13 million versus 2.57 million).
On Tiny-ImageNet, we also achieved comparable results to the Joint-SNN model in 4 time steps, demonstrating that CogniSNN performs well when facing large static datasets.
@@
Figure 3: Comparison of results of OR and other operations on DVS-Gesture with ER-RGA-7 based CogniSNN.
5.4 Depth-Scalability
The depth-scalability of CogniSNN is heavily based on the usage of skip connections within ResNodes. As described in Section 2.1, early research is generally based on the following four configurations: 1) without skip connection, 2) ADD (addition), 3) AND, and 4) IAND. We compare our skip connection (OR gate) with the four aforementioned configurations based on the ER-driven CogniSNN. Furthermore, all models with/without configuring different skip connection mechanisms are trained on the DVS-Gesture.
@@ 5.4 Depth-Scalability
[…]
5.5 Path-Plasticity
We apply the critical path-based LwF with $K=1$ on the WS-driven CogniSNN configured with 32 ResNodes. Comparable results are observed with the ER-driven CogniSNN and are provided in Section 6 of Supplementary Material. Tables 4 and 5 present the results, maintaining the significance of each column consistent with the corresponding column in Tables 2 and 3. The first row presents results from the application of the fine-tuned CogniSNN using vanilla LwF on both tasks. The subsequent two rows display results utilizing the proposed critical path-based LwF, implemented based on paths with the highest and lowest betweenness centrality, respectively.
Table 4: Comparison of accuracy ($\%$) between similar tasks.
Method CIFAR100 CIFAR10 CIFAR100